The operational plan for Sites 980 and 981 (FENI-1 and FENI-2) called for coring to approximately 110 mbsf and 225 mbsf, respectively, for three APC holes at each site. The vessel left Leith the evening of 7 July. At midnight on 9 July the ship slowed down to 6.0 kt. The seismic gear was deployed and tested for 1.5 hr, and approximately 6 hr were spent surveying these first two coring sites near the crest of the Feni sediment drift. Based on the survey results, we located Site 980 (FENI-1) about 0.5 nm south of the originally proposed site in order to better target an expanded upper seismic sequence (Pleistocene). Site 981 (FENI-2) was located about 0.5 nautical miles (nmi) northeast of the original site in the Leg 162 Prospectus and targeted an expanded lower seismic section (Pliocene). After the site survey was completed, the ship returned to Site 980 and dropped the beacon at 0846 hours on 10 July 95.
A standard APC/XCB bottom-hole assembly (BHA) was used for all holes at Sites 980 and 981, including a nonmagnetic drill collar. At both sites, subsequent holes (B and C) were offset 15 m to the north. The mudline was established for each hole. The APC firing depth was offset by a few meters for subsequent holes to establish continuous sediment sections.
Coring at Site 980 proceeded smoothly, with only one minor incident. The first APC shot at Hole 980B from 2179 mbrf, which was 3 m higher than for Hole 980A, recovered a full core that was discarded. The pipe was raised to 2174 mbrf, or a full 8 m higher than where Hole 980A was spudded from, and recovered 4.17 m of core. Target depth was reached in each hole. Core orientation was conducted using the Tensor tool on Cores 980B-3H through 13H, and Cores 980C-4H through 14H. Operations at Site 980 were terminated when the beacon was recovered at 2300 hrs on 11 July.
Site 981 (FENI-2) is approximately 2 nmi southeast of Site 980, so the vessel was maneuvered in DP mode to the global positioning system (GPS) coordinates defined by the site survey. This took about 2.4 hr. The drill string was pulled to a depth of 2059.1 m or approximately 120 m above the seafloor during this transit. At 0120 hrs on 12 July the positioning beacon was dropped. Coring proceeded without incident with two minor exceptions. First, an incomplete stroke occurring on Core 981A-32H unseated the oil saver when the APC barrel was rigged for Core 981A-34H, at which point the sinker bars were laid out and operations were switched to the forward coring line. Second, coring was suspended in Hole 981A when a maximum overpull of 50,000 pounds occurred on Core 981A-34H. Scientific target depth was reached in all holes.
Core orientation was conducted using the Tensor tool on Cores 981A-3H through 34H and Cores 981C-4H through 14H. In addition, 15 temperature measurements were made using the ADARA tool. Eleven consecutive temperature measurements were made on Cores 981C-4H through 14H, and additional measurements were made on Cores 981C-16H, 18H, 20H, and 22H. Core 981C-22H required washover of 5.0 m and an overpull of 100,000 lb prior to retrieval. Temperature measurements were suspended at that point. Occupation of Site 981 ended when the vessel was secured for transit and got underway at 0315 hr on 15 July 1995.
To Operations Site 982
162 Table of Contents